OpenStudy (anonymous):

Which function below is the inverse of f(x) = The quantity of four x minus three, over two.? f−1(x) = two over the quantity of four x minus three f−1(x) = two over the quantity of three minus four x f−1(x) = The quantity of two x minus three, over four f−1(x) = The quantity of two x plus three, over four

OpenStudy (anonymous):

f-1(x)=2x+3/4

OpenStudy (ranga):

\[f(x) = \frac{ 4x - 3 }{ 2 }\quad f^{-1}(x) = ?\]

OpenStudy (ranga):

Write it as:\[y = \frac{ 4x - 3 }{ 2 }\] Replace x with y and y with x and then solve for y.

OpenStudy (ranga):

That will be your inverse function.

OpenStudy (ranga):

y = (4x - 3) / 2 interchange x and y x = (4y - 3) / 2 multiply both sides by 2 2x = 4y - 3 add 3 to both sides 2x + 3 = 4y divide both sides by 4 (2x + 3) / 4 = y So the inverse function = (2x + 3) / 4